OpenStudy (anonymous):

Please Help!!! Find the interest due on $1,200 at 8% for 240 days. 63.33 64.00 63.67

OpenStudy (anonymous):

SI = PTR /100 P = $1200 given T = 240 Days given and R = 8/100 = 0.08 Given substitute them and u get the answer